OpenStudy (anonymous):

Write a formula for the sequence: 10, 17, 24, 31,... starting at n=1.

OpenStudy (anonymous):

31-24=7 24-17=7 17-10=7 Seeing a pattern yet?

OpenStudy (anonymous):

here a =10 d=17-10=7 since tn =a+(n-1)d =10+(n-1)7 =10+7n-7 =7n+3 tn =7n+3 which is the required formula Now at n=1 we have \[t_1=7 \times 1+3 =7+3=10\]